OpenStudy (anonymous):

will fave and give medal! please and thanks :) Problem Solving. Show solutions: a) The difference between the measures of an angle and its supplement is forty eight. Find the measure of both angles. b) The ratio of the interior angles of a Euclidean triangle is 3:4:5. Find the measure of each angle. c) The length of a rectangle is three more than twice its width. What are the dimensions of a rectangle if its perimeter is 96 cm.

OpenStudy (dmb1215):

a i would say is 52

OpenStudy (dmb1215):

what is a eeuclidean triangle

OpenStudy (anonymous):

180-48=132 3+4+5=12 180/12= 15 3 x 15 = 45 4 x 15 = 60 5 x 15= 75 45,60,75=180

OpenStudy (anonymous):

45:60:75

OpenStudy (anonymous):

i got answer of ur b one b: first sum of three angles is 12 i mean 3+4+5 = 12 then u do \[\frac{ 3 }{ 4 }\times180=45\] and then same u go with\[\frac{ 4 }{ 12 }\times180=60\]and\[\frac{ 5 }{ 12 }\times180=75\]here u go u got 45,60 and 75 as answer